### v1

@@ 4 Discussion
A comparison of the fragmentation chronology with the photometric evolution reveals a systematic delay between the physical breakup of the nucleus and the subsequent rise in brightness. Back–extrapolation of the I–II separation places the primary fragmentation event around 2025 Nov. 1 (within a few days), whereas the major outburst detected in the LOOK and ZTF photometry occurred between Nov. 2.54 and Nov. 4.51 UTC. The delay of roughly 1133 days suggests that freshly exposed interior ices remained extremely cold immediately after fragmentation and required time for a thermal wavae to propagate inward before vigorous sublimation and dust release could commence. The secondary fragmentation of II shows a similar pattern: the IIa–IIb split occurred between Nov. 8.56 and Nov. 9.82, yet IIb brightened substantially only by the final HST epoch on Nov. 10.53, again implying a $\sim$1 day delay. The smaller ($\sim$0.2 mag) outburst detected shortly before the Nov. 10 LCO observations likewise fits within this sequence of delayed activation. These lags are naturally explained through basic thermal physics. For porous cometary material with thermal diffusivity $\kappa\sim 10^{-7}$$10^{-8}~\mathrm{m^{2}\,s^{-1}}$ (Davidsson and Rickman, 2014), the penetration depth of the thermal wave is $\delta\sim\sqrt{\kappa t}$. Reaching temperatures of $\sim$180–200 K, sufficient for rapid $\mathrm{H_{2}O}$ sublimation at $r_{h}\approx 0.9$ au, requires warming layers several centimeters deep. This corresponds to heating timescales of $t\sim 1$33 days, in excellent agreement with the observed photometric delays. Thus, the newly exposed surfaces created by fragmentation require several rotation cycles before effective sublimation can drive strong dust production.
@@
The arclets surrounding fragment A reinforce this interpretation. Arclet-like features have been documented in several comets (e.g., Harris et al., 1997; Boehnhardt, 2004), but in C/2025 K1 they are observed only in the first HST epoch (Nov. 8.56) and are absent in the subsequent images. Their fixed curvature and lack of outward motion over the paired frames within that epoch indicate short-lived, confined structures rather than propagating waves. This behavior differs markedly from the large-scale, quasi-periodic pressure waves proposed by Farnham et al. (2001) for comet C/1996 B2 (Hyakutake), where arclets appeared between the primary nucleus and a companion fragment and propagated coherently across the coma. In C/2025 K1, the arclets are anchored directly on fragment I, do not move outward, and show no evidence for global coherence.
@@
Harris et al. (1997) suggested that the arclets in Hyakutake were dominated by gas emission; within the STIS F28X50LP bandpass this would most plausibly arise from NH2 $\tilde{A}^{\,2}A_{1}\;-\;\tilde{X}^{\,2}B_{1}$. Recent spectra of C/2025 K1 likewise show strong NH2 emission and very weak CN, C2, and C3 (Ganesh et al., 2025). However, several lines of evidence argue against a gas–emission origin for the arclets in C/2025 K1. First, the features are sharply delineated, geometrically thin, morphologies inconsistent with the smoother, more diffuse structures expected from NH2 fluorescence, whose parent NH3 has short photochemical lifetimes at $r_{h}\sim 0.9$ au. Second, the arclets are tightly centered on fragment A rather than appearing between fragments, as in Hyakutake. Third, the lack of outward expansion is incompatible with gas moving at $\sim$0.5–1 km s-1. Together, these properties strongly favor an interpretation in which the arclets are thin dust shells released when freshly exposed surfaces on fragment I were first illuminated and began to warm, an origin fully consistent with the thermal-delay scenario inferred from the lightcurve.
@@
If the observed outbursts are instead linked to rotational instabilities, simple models allow order-of-magnitude estimates of fragment sizes. Outgassing torques can spin up fragments to failure in 1–3 days for typical cometary densities ($\rho\sim 500~\mathrm{kg\,m^{-3}}$) and cohesive strengths ($S\sim 10$–100 N m-2) (Jewitt et al., 2020). Using the observed delays between inferred breakup times and increased activity, fragments undergoing secondary disruption (such as B1 and B2) are consistent with radii of roughly 5–40 m, assuming active fractions near unity. These dimensions comfortably match the slow, near–escape-velocity separation of IIa and IIb.
@@ 5 Summary
The combined HST and LOOK observations capture C/2025 K1 (ATLAS) only days after a major breakup, providing one of the earliest and clearest views of an evolving cometary fragmentation sequence. HST resolves at least five fragments and reveals a hierarchical disruption, including a slow secondary split of fragment II. Ground-based photometry contextualizes these events within the broader activity evolution, linking the post-perihelion 0.9 mag outburst and subsequent brightening episodes to the emergence and activation of individual fragments. Fragment motions imply that the primary breakup separated kilometer-scale pieces at a few meters per second, while the IIa–IIb split proceeded at near–escape velocity. A consistent 1–3 day delay between physical fragmentation and enhanced dust production reveals how newly exposed subsurface layers warm and begin to release volatiles. The thin, short-lived arclets around fragment I in the first HST epoch demonstrate that the initial exposure produces only weak, surface-confined dust release, followed by stronger activity as heat diffuses into deeper, more volatile-rich layers. These patterns indicate that the volatile inventory is vertically and laterally structured: surface layers are devolatilized, while interior regions retain more abundant or more easily mobilized volatiles. The absence of immediate, explosive activity argues against hypervolatile-driven disruption and instead suggests that sublimation is throttled by the time required for vapor pressure to build within porous interior pathways.
The morphology, kinematics, and activation behavior therefore point to a nucleus-wide mechanical failure that exposed heterogeneously mixed interior ices, followed by thermally mediated volatile release governed by diffusion and pore connectivity. This behavior parallels, but greatly scales up, the delayed activation seen after cliff collapses on 67P, where the timing and morphology of dust release reflect how heat propagates, how trapped volatiles are stored, and how efficiently material is transported to the surface. These observations provide rare, time-resolved constraints on the mixing, storage, and release of volatiles in a dynamically new comet near perihelion. They show that the early evolution of a disrupted nucleus records both the mechanical architecture of the breakup and the thermal–physical response of newly exposed interior layers, offering direct insight into volatile stratification, permeability, and the efficiency of dust–gas coupling in cometary interiors.
@@ Acknowledgements
This work makes use of observations from the Las Cumbres Observatory global telescope network. Observations with the LCOGT 1 m were obtained as part of the LCO Outbursting Objects Key (LOOK) Project (Proposal IDs: LTP2025A-004). This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with program ID 18135 (PI D. Bodewits). We gratefully acknowledge the outstanding support of the HST/STScI scheduling and operations staff during the execution of Program 18135. In particular, we thank Alison Vick, Sten Hasselquist, and TalaWanda Monroe at the Space Telescope Science Institute for their expert assistance in obtaining these challenging observations. We would also express our gratitude to Qicheng Zhang for reporting his observations of the October outburst and fragmentation to the MPC, which helped us with the interpretation of the sequence of events. Supporting observations were obtained by the Comet Chasers education and outreach program (https://www.cometchasers.org/ ), led by HU, which is funded by the UK Science and Technology Facilities Council (via the DeepSpace2DeepImpact Project), the Open University and Cardiff University. It accesses the 2m, 1m and 0.4m LCOGT telescopes through the Schools Observatory/Faulkes Telescope Project (TSO2025B-01 The Schools’ Observatory/DFET), which is partly funded by the Dill Faulkes Educational Trust, and through the LCO Global Sky Partners Programme (LCOEPO2023B-013). Observations were made by the following schools and clubs: Institut d’Alcarràs, Catalonia, Spain; Jelkovec High School, Zagreb, Croatia; Louis Cruis Astronomy Club, Brazil; St Marys Catholic Primary School, Bridgend, UK; The Coopers Company & Coborn School, Upminster, UK (as part of the BAA Work Experience Project).
We acknowledge with thanks the comet observations from the COBS Comet Observation Database contributed by observers worldwide and used in this research.

### v2

@@ 4 Discussion
A comparison of the fragmentation chronology with the photometric evolution reveals a systematic delay between the physical breakup of the nucleus and the subsequent rise in brightness. Back–extrapolation of the I–II separation places the primary fragmentation event around 2025 Nov. 1 (within a few days), whereas the major outburst detected in the LOOK and ZTF photometry occurred between Nov. 2.54 and Nov. 4.51 UTC. The delay of roughly 1133 days suggests that freshly exposed interior ices remained extremely cold immediately after fragmentation and required time for a thermal wave to propagate inward before vigorous sublimation and dust release could commence. The secondary fragmentation of II shows a similar pattern: the IIa–IIb split occurred between Nov. 8.56 and Nov. 9.82, yet IIb brightened substantially only by the final HST epoch on Nov. 10.53, again implying a $\sim$1 day delay. The smaller ($\sim$0.2 mag) outburst detected shortly before the Nov. 10 LCO observations likewise fits within this sequence of delayed activation. These lags are naturally explained through basic thermal physics. For porous cometary material with thermal diffusivity $\kappa\sim 10^{-7}$$10^{-8}~\mathrm{m^{2}\,s^{-1}}$ (Davidsson and Rickman, 2014), the penetration depth of the thermal wave is $\delta\sim\sqrt{\kappa t}$. Reaching temperatures of $\sim$180–200 K, sufficient for rapid $\mathrm{H_{2}O}$ sublimation at $r_{h}\approx 0.9$ au, requires warming layers several centimeters deep. This corresponds to heating timescales of $t\sim 1$33 days, in excellent agreement with the observed photometric delays. Thus, the newly exposed surfaces created by fragmentation require several rotation cycles before effective sublimation can drive strong dust production.
An alternative, but physically consistent, interpretation of the observed delay is that it reflects the time required to re-establish a mechanically unstable dust mantle on freshly exposed fragment surfaces, rather than solely the time needed to heat subsurface ice to sublimation temperatures. Immediately after fragmentation, the fragment surfaces are expected to expose an intimate mixture of ice and dust representative of the bulk nucleus. Upon solar illumination, such surfaces can begin vigorous gas sublimation on short timescales (hours), but without substantial dust release. Significant dust production may only commence once a porous dust mantle of sufficient thickness has re-formed and thermally decoupled from the underlying sublimation front.
Order-of-magnitude energy-balance considerations support this scenario. For heliocentric distances near $r_{h}\approx 0.9$ au, the net absorbed solar flux is sufficient to heat newly exposed ice–dust mixtures from $\sim$100 K to water-sublimation temperatures ($\sim$200 K) in only a few hours. However, most of the absorbed energy over the following days is expended on sublimating near-surface water ice, allowing a dust mantle to grow to a thickness of a few centimeters over $\sim$1–3 days. Thermophysical models of comet 67P show that mantles of comparable thickness naturally develop under similar conditions (Davidsson et al., 2022) and that such mantles are prone to strong thermal gradients and mechanical failure once radiative equilibrium is established at the surface (Alí-Lagoa et al., 2015; Attree et al., 2018; Davidsson et al., 2022).
An additional and complementary consequence of this delayed thermal activation is outgassing-driven spin-up of the newly exposed fragments. As sublimation intensifies several days after perihelion, asymmetric gas emission can exert substantial torques on irregularly shaped nuclei, driving rapid changes in rotation state. Rotational instability is therefore an efficient mechanism for both continued fragmentation and enhanced dust release following the initial breakup. Sub-kilometer cometary nuclei are especially vulnerable to this process, as even modest outgassing can spin them up to structural failure on timescales of days (Samarasinha and Mueller, 2013; Jewitt, 2021). A well-known example is the small nucleus of 289P/Blanpain (radius $\sim$160 m), which shows only weak present-day activity yet is widely interpreted as the remnant of a rotationally driven disintegration event that produced the Phoenicid meteoroid stream (Jewitt, 2006; Kasuga and Jewitt, 2019; Kasuga, 2025).
For C/2025 K1, the fragment cross sections measured by HST provide only upper limits on the true nucleus sizes due to coma contamination, implying that the actual solid fragments are likely significantly smaller than the inferred kilometer-scale values. The observed near–escape-velocity separations of the fragments (0.5–10 m s-1), and in particular the very low relative speed of IIa and IIb ($\sim$0.5 m s-1), are consistent with rotational disruption rather than direct gas-drag ejection. These velocities are far below typical gas expansion speeds ($\sim$500 m s-1 for H2O at 1 au) and are difficult to reconcile with classical gas-coupling models (Whipple, 1951), which predict sub–m s-1 ejection speeds only for meter-scale debris even at the perihelion distance of C/2025 K1. Rotational instability therefore provides a physically plausible pathway linking delayed heating, fragment spin-up, secondary breakup, and the observed dust production. In this framework, delayed heating not only regulates the onset of sublimation but also governs when rotational stresses peak, naturally coupling the timing of enhanced activity to subsequent fragmentation.
@@
The arclets surrounding fragment I reinforce this interpretation. Arclet-like features have been documented in several comets (e.g., Harris et al., 1997; Boehnhardt, 2004), but in C/2025 K1 they are observed only in the first HST epoch (Nov. 8.56) and are absent in the subsequent images. Their fixed curvature and lack of outward motion over the paired frames within that epoch indicate short-lived, confined structures rather than propagating waves. This behavior differs markedly from the large-scale, quasi-periodic pressure waves proposed by Farnham et al. (2001) for comet C/1996 B2 (Hyakutake), where arclets appeared between the primary nucleus and a companion fragment and propagated coherently across the coma. In C/2025 K1, the arclets are anchored directly on fragment I, do not move outward, and show no evidence for global coherence.
@@
Harris et al. (1997) suggested that the arclets in Hyakutake were dominated by gas emission; within the STIS F28X50LP bandpass this would most plausibly arise from NH2 ($\tilde{A}^{\,2}A_{1}\;-\;\tilde{X}^{\,2}B_{1}$). Recent spectra of C/2025 K1 likewise show strong NH2 emission and very weak CN, C2, and C3 (Ganesh et al., 2025). However, several lines of evidence argue against a gas–emission origin for the arclets in C/2025 K1. First, the features are sharply delineated, geometrically thin, morphologies inconsistent with the smoother, more diffuse structures expected from NH2 fluorescence, whose parent NH3 has short photochemical lifetimes at $r_{h}\sim 0.9$ au. Second, the arclets are tightly centered on fragment I rather than appearing between fragments, as in Hyakutake. Third, the lack of outward expansion is incompatible with gas moving at $\sim$0.5–1 km s-1. Together, these properties strongly favor an interpretation in which the arclets are thin dust shells released when freshly exposed surfaces on fragment I were first illuminated and began to warm, an origin fully consistent with the thermal-delay scenario inferred from the lightcurve.
An important implication of this interpretation concerns the relationship between observed coma abundance ratios and the true bulk composition of the nucleus. Prior to fragmentation, strong near-surface stratification is expected, with dust-rich layers overlying deeper sublimation fronts for H2O, CO2, and CO. In such cases, coma abundances reflect the combined effects of temperature gradients, diffusion, and recondensation rather than the intrinsic ice abundances at depth (Marboeuf and Schmitt, 2014; Davidsson et al., 2022). Immediately after fragmentation, however, all volatile species are temporarily exposed at the surface at their bulk nucleus abundances. The observations presented here empirically constrain this window to the first $\sim$1–3 days following breakup. Any gas abundance ratios measured during this brief interval would therefore provide an unusually direct probe of nucleus interior composition. Moreover, comparisons between fragments during this early phase would offer a powerful test of whether cometary nuclei are compositionally homogeneous or heterogeneous at depth. We note that both COS and STIS spectroscopic observations that can address this were taken as part of GO-18135 and will be published separately.
@@
If the observed outbursts are instead linked to rotational instabilities, simple models allow order-of-magnitude estimates of fragment sizes. Outgassing torques can spin up fragments to failure in 1–3 days for typical cometary densities ($\rho\sim 500~\mathrm{kg\,m^{-3}}$) and cohesive strengths ($S\sim 10$–100 N m-2) (Jewitt et al., 2020). Using the observed delays between inferred breakup times and increased activity, fragments undergoing secondary disruption (such as II) are consistent with radii of roughly 5–40 m, assuming active fractions near unity. These dimensions comfortably match the slow, near–escape-velocity separation of IIa and IIb.
@@ 5 Summary
The combined HST and LOOK observations capture C/2025 K1 (ATLAS) only days after a major breakup, providing one of the earliest and clearest views of an evolving cometary fragmentation sequence. HST resolves at least five fragments and reveals a hierarchical disruption, including a slow secondary split of fragment II. Ground-based photometry contextualizes these events within the broader activity evolution, linking the post-perihelion 0.9 mag outburst and subsequent brightening episodes to the emergence and activation of individual fragments. Fragment motions imply that the primary breakup separated kilometer-scale pieces at a few meters per second, while the IIa–IIb split proceeded at near–escape velocity. A consistent 1–3 day delay between physical fragmentation and enhanced dust production reveals how newly exposed subsurface layers warm and begin to release volatiles. The thin, short-lived arclets around fragment I in the first HST epoch demonstrate that the initial exposure produces only weak, surface-confined dust release, followed by stronger activity as heat diffuses into deeper, more volatile-rich layers. These patterns indicate that the volatile inventory is vertically and laterally structured: surface layers are devolatilized, while interior regions retain more abundant or more easily mobilized volatiles. The absence of immediate, explosive activity argues against hypervolatile-driven disruption and instead suggests that sublimation is throttled by the time required for vapor pressure to build within porous interior pathways. The empirically constrained 1–3 day window between fragmentation and dust activation therefore implies a brief interval during which coma gas abundances may closely reflect bulk nucleus composition, underscoring the importance of rapid compositional measurements following breakup events.
The morphology, kinematics, and activation behavior therefore point to a nucleus-wide mechanical failure that exposed heterogeneously mixed interior ices, followed by thermally mediated volatile release governed by diffusion and pore connectivity and, in some fragments, by outgassing-driven rotational instability. This behavior parallels, but greatly scales up, the delayed activation seen after cliff collapses on 67P, where the timing and morphology of dust release reflect how heat propagates, how trapped volatiles are stored, and how efficiently material is transported to the surface. These observations provide rare, time-resolved constraints on the mixing, storage, and release of volatiles in a dynamically new comet near perihelion. They show that the early evolution of a disrupted nucleus records both the mechanical architecture of the breakup and the thermal–physical response of newly exposed interior layers, offering direct insight into volatile stratification, permeability, and the efficiency of dust–gas coupling in cometary interiors.
@@ Acknowledgements
This work makes use of observations from the Las Cumbres Observatory global telescope network. Observations with the LCOGT 1 m were obtained as part of the LCO Outbursting Objects Key (LOOK) Project (Proposal IDs: LTP2025A-004). This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with program ID 18135 (PI D. Bodewits). We gratefully acknowledge the outstanding support of the HST/STScI scheduling and operations staff during the execution of Program 18135. In particular, we thank Alison Vick, Sten Hasselquist, and TalaWanda Monroe at the Space Telescope Science Institute for their expert assistance in obtaining these challenging observations. We would also express our gratitude to Qicheng Zhang for reporting his observations of the October outburst and fragmentation to the MPC, which helped us with the interpretation of the sequence of events. Supporting observations were obtained by the Comet Chasers education and outreach program (https://www.cometchasers.org/ ), led by HU, which is funded by the UK Science and Technology Facilities Council (via the DeepSpace2DeepImpact Project), the Open University and Cardiff University. It accesses the 2m, 1m and 0.4m LCOGT telescopes through the Schools Observatory/Faulkes Telescope Project (TSO2025B-01 The Schools’ Observatory/DFET), which is partly funded by the Dill Faulkes Educational Trust, and through the LCO Global Sky Partners Programme (LCOEPO2023B-013). Observations were made by the following schools and clubs: Institut d’Alcarras, Catalonia, Spain; Jelkovec High School, Zagreb, Croatia; Louis Cruis Astronomy Club, Brazil; St Marys Catholic Primary School, Bridgend, UK; The Coopers Company & Coborn School, Upminster, UK (as part of the BAA Work Experience Project). Parts of this research were carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. We acknowledge with thanks the comet observations from the COBS Comet Observation Database contributed by observers worldwide and used in this research.